OpenStudy (math_bender_15):

Karen runs 5 miles in 58 minutes. At the same rate, how many miles would she run in 87 minutes?

hero (hero):

Hint: Set up the following proportion \(\large\frac{5 \text{miles}}{x \text{miles}} = \frac{58 \text{minutes}}{87 \text{minutes}} \)

hero (hero):

Then after canceling miles and minutes you're simply left with: \(\large\frac{5}{x} = \frac{58}{87}\) Now you can just cross multiply and solve for x.

OpenStudy (math_bender_15):

7.5

hero (hero):

x = 7.5 Good job

OpenStudy (math_bender_15):

thanks